Suppose that N= sqrt X + 2 models the number of cases of an infection, in millions, of a disease x years from now.
a) Approximately how many cases of the infection will there be 16 years from now?
B. In approximately how many years will there be 7 million cases?

Answer by oscargut(2103)


a) if x=16 then N= sqrt(16)+2=6 so the response is 6 cases
b) if N=7 then 7=sgrt(x)+2 then x is 25 so the response is 25 years